OpenStudy (annalee607):

It costs a family $324.00 to buy a ten-ft-by-12-ft rug. If the cost is proportional to the area, what will it cost the family to buy a 15-ft-by-18-ft rug?

myininaya (myininaya):

Do you know what it means for the cost to be proportional to the area?

OpenStudy (annalee607):

no

myininaya (myininaya):

This means there is some constant, let's call it, k such that cost=k*area

myininaya (myininaya):

ok now we have an equation do you know how to find the area of a 10 ft by 12 ft rug?

OpenStudy (annalee607):

multiply them?

myininaya (myininaya):

right

myininaya (myininaya):

so the area=10*12 ft^2 when we have the cost is 324 this is a point on our line cost=k*area use this point (area=10*12, cost=324) to find k

myininaya (myininaya):

solve for k and I will check your k then we will move on to the final question of this problem

OpenStudy (annalee607):

I am very confused

myininaya (myininaya):

we were given cost=k*area when it told us that the cost was proportional to the area we were also given when the area=10*12 we have the cost=324 this is enough information to find our constant of proportionally (aka k)

myininaya (myininaya):

you need to enter area=10*12 and cost=324 into cost=k*area to find k do you still not understand what I'm asking you to do? please let me know.

OpenStudy (annalee607):

324=k*120

myininaya (myininaya):

right solve for k

OpenStudy (annalee607):

324/120=2.7

myininaya (myininaya):

cool so you just found the constant of proportionality

myininaya (myininaya):

so we can find the cost for any area rug or find the area of the rug given the cost

myininaya (myininaya):

so we have cost=2.7*area

myininaya (myininaya):

what will it cost the family to buy a 15-ft-by-18-ft rug this is your question

myininaya (myininaya):

you are given the area basically and you are asked to find the cost using the equation cost=2.7*area

OpenStudy (annalee607):

cost=2.7*270 cCost=$729.00

myininaya (myininaya):

beautiful

myininaya (myininaya):

you want me to show a slightly different way you could have done this problem ?

myininaya (myininaya):

you could have just had one equation instead of even bothering with the constant of proportionality that is you could have just solved \[\frac{\text{ area of second rug }}{\text{ area of first rug }}=\frac{\text{ cost of second rug }}{\text{ cost of first rug }} \\ \frac{x}{324}=\frac{270}{120} \\ \text{ multiply both sides by } 324 \\ x=\frac{270}{120} \cdot 324=729\]

myininaya (myininaya):

if you didn't like the first way please read that last post

myininaya (myininaya):

lol I put the costs on the opposite side but whatever a=b means b=a :p